# MIRA: Medical Time Series Foundation Model for Real-World Health Data

### v2

@@ 4.5 Ablation Analysis
Figure 3: (Left) Training curves on Medical corpora of different MIRA model sizes. (Right) Gating scores for experts across different layers in the there different frequence datasets.
@@
Expert Activation Analysis. We further examine the trade-off between predictive performance and inference efficiency by varying the number of activated experts ($top_{k}$), as shown in the right-hand side of Table 5. Activating two experts ($top_{2}$) achieves the best balance, reaching the lowest RMSE of 0.154 with an inference speed of 0.101 s/iter. Increasing to four experts provides no meaningful improvement in performance but increases inference time to 0.112 s/iter. Further increasing to six or eight experts leads to slower inference with diminishing returns. Conversely, using a single expert ($top_{1}$) sacrifices accuracy (0.160) for only a marginal speed gain. These findings indicate that using two experts offers an optimal trade-off between efficiency and accuracy for scalable deployment.
@@ Appendix B Time Quantization Procedure
The procedure begins with a predefined initial resolution (e.g., 1.0) and iteratively reduces it by a shrink factor (e.g., 10) until all quantized time points become unique. In each iteration, the algorithm rounds the original timestamps to the current resolution and checks for duplicates. If uniqueness is achieved, the quantized sequence is returned.
If, after a fixed number of iterations, uniqueness cannot be guaranteed, a fallback mechanism is triggered. In this case, small additive jitter is applied to break ties while preserving temporal ordering. The jitter is dynamically determined based on the magnitude of the input values, ensuring that the perturbation is both numerically safe and practically negligible.
Finally, the adjusted timestamps are cast to 32-bit floating-point precision and verified again for uniqueness. If rare collisions still exist due to floating-point rounding, a second corrective jitter step is applied.
This method ensures that all quantized timestamps are:
  • Strictly increasing, preserving temporal semantics;
  • Resolution-adaptive, automatically selecting the finest possible scale without manual tuning;
  • Numerically stable, avoiding floating-point collisions and degenerate attention behavior.
@@ C.1 Theoretical Properties
Linear Angle Scaling
@@
Proof. Direct from Equation (1) in the main text. The linear formulation preserves temporal causality while maintaining bounded rotation magnitudes through exponentially decaying frequencies:
$\omega_{i}=\exp\left(-\frac{2i}{d}\ln 10000\right)$ (14)
This ensures higher dimensions receive progressively smaller rotations.
Relative Position Encoding
The inner product between rotated vectors depends only on the temporal difference:
$\langle R_{\Theta,t_{i}}q,R_{\Theta,t_{j}}k\rangle=\langle q,R_{\Theta,|t_{i}-t_{j}|}k\rangle$ (15)
Proof. Using rotation matrix properties:
$R_{\Theta,t_{i}}^{\top}R_{\Theta,t_{j}}$ $=\prod_{k=0}^{d/2-1}\begin{bmatrix}\cos(\omega_{k}(t_{j}-t_{i}))&\sin(\omega_{k}(t_{j}-t_{i}))\\ -\sin(\omega_{k}(t_{j}-t_{i}))&\cos(\omega_{k}(t_{j}-t_{i}))\end{bmatrix}$ (16)
$=R_{\Theta,t_{j}-t_{i}}=R_{\Theta,|t_{i}-t_{j}|}$ (17)
The absolute difference emerges from $\cos(-\theta)=\cos(\theta)$ and $\sin(-\theta)=-\sin(\theta)$.
C.2 Rotation Matrix Construction
For a $d$-dimensional model (even dimension), define block-diagonal rotation matrix:
$R_{\Theta,t}=\bigoplus_{i=0}^{d/2-1}\begin{bmatrix}\cos(\omega_{i}t)&-\sin(\omega_{i}t)\\ \sin(\omega_{i}t)&\cos(\omega_{i}t)\end{bmatrix}$ (18)
with key properties:
  • Norm Preservation: $\|R_{\Theta,t}x\|=\|x\|,\ \forall x\in\mathbb{R}^{d}$
  • Temporal Ordering: $t_{1}<t_{2}\Rightarrow\theta_{i}(t_{1})<\theta_{i}(t_{2})$
  • Differentiable: $\frac{\partial R_{\Theta,t}}{\partial t}$ exists $\forall t>0$
@@ Appendix D Neural ODE
Gradient Computation via Adjoint Sensitivity
For end-to-end training, gradients of a loss function $\mathcal{L}$ (computed based on the final prediction) with respect to the ODE parameters $\theta_{ODE}$ and the initial state $h_{N}$ are efficiently computed using the adjoint sensitivity method [39]. This involves solving a backward-in-time ODE for the adjoint state $a(s)=\partial\mathcal{L}/\partial h(s)$:
$\frac{da(s)}{ds}=-a(s)^{\top}\frac{\partial f_{ODE}}{\partial h}(s-t_{N},h(s);\theta_{ODE})$ (20)
with the terminal condition $a(t_{N+1})=\partial\mathcal{L}/\partial h^{\prime}_{N+1}$. The required gradients are then computed via integrals involving the adjoint state $a(s)$.
Stability Considerations
To promote numerical stability and ensure the dynamics function $f_{ODE}$ remains well-behaved (e.g., satisfies Lipschitz conditions), techniques like spectral normalization can be applied to the weights $\theta_{ODE}$ of the neural network implementing $f_{ODE}$. This helps control the Lipschitz constant $L$ of $f_{ODE}$ with respect to $h$, contributing to bounded error propagation during integration as characterized by bounds of the form:
$\|h_{ODE}(t_{N+1})-h_{\text{true}}(t_{N+1})\|\leq\mathcal{O}\left(\frac{M}{L}(e^{L(t_{N+1}-t_{N})}-1)\right)$ (21)
where $M$ relates to the local truncation error of the numerical solver.
@@
0:  Hidden state $\mathbf{h}_{N}$, timestamps $t_{N}$, $t_{N+1}$
0:  Evolved state $\mathbf{\tilde{h}}(t_{N+1})$
@@
2:  Solve ODE: $\mathbf{\tilde{h}}(t_{N+1})\leftarrow\texttt{odeint}(f_{\text{ODE}},\mathbf{h}_{N},[t_{N},t_{N+1}])$
3:  Return final state: $\mathbf{\tilde{h}}(t_{N+1})[-1]$
D.1 Theoretical Guarantees
Our design satisfies the Picard-Lindelöf theorem conditions:
Theorem 1.
Let $f_{\text{ODE}}$ be spectrally normalized with maximum singular value $\sigma_{\max}$. Then for prediction horizon $\Delta t=t_{N+1}-t_{N}$, the state evolution admits a unique solution satisfying:
$\|\mathbf{\tilde{h}}(t_{N+1})-\mathbf{\tilde{h}}_{N}\|\leq\sigma_{\max}\Delta te^{\sigma_{\max}\Delta t}$ (22)
This bounded evolution property ensures numerical stability during long-horizon predictions.
@@ Appendix F Model Configurations and Training Detail
Informed by recent scaling studies [28], we design MIRA in three model scales: $MIRA_{small}$ (73M parameters), $MIRA_{base}$ (114M), and $MIRA_{large}$ (455M), enabling flexible deployment across different hardware constraints. All models are trained on up to eight NVIDIA 80GB A100 GPUs with a micro-batch size of 128 and a maximum sequence length of 512. We pre-trained one epoch with each training step processes approximately 65,000 time points. We consider forecast horizons of ${24,32,48,64}$ for short-term and long-term evaluation. Following standard practice, we apply an auxiliary load balancing loss with weight $\alpha=0.02$ to encourage expert utilization.
@@ Appendix G Downstream Tasks and Datasets
Table 8: Characteristics of the WHO FluNet Dataset.
Dataset Time Points Missing Rate Frequency Irregular Sampling Signal Type
WHO FluNet 5.27M 61.52% Weekly Case Counts
CINC 2012 5.28M 0% Variable-specific ECG, Blood Pressure (ABP/NIBP), Respiration Rate, SpO2, Laboratory Measurements, etc.
MIT-BIH 31.1M 0% 360 Hz ECG
CinC 2019 39.5M 68.28% Hourly Heart Rate, Respiration Rate, Temperature, Blood Pressure
Johns Hopkins COVID-19 Dataset 3.81M 0% Daily Case Counts, Death Counts
CDC-Influenza Hospitalizations Admissions 3.07M 27.1% Weekly Hospitalized Case Counts
@@
WHO FluNet [^5] https://www.who.int/tools/flunet These epidemiological datasets from the World Health Organization comprise country-level weekly flu case reports (FluNet).
@@ Appendix H Baselines
Sundial [31] introduces a family of scalable time series foundation models trained on the extensive TimeBench dataset, comprising one trillion time points. Utilizing the novel TimeFlow Loss, Sundial enables generative pretraining of Transformers without discrete tokenization, achieving state-of-the-art zero-shot performance across diverse forecasting benchmarks.
@@ Appendix I Evaluation Metrics
Models Lag-llama TimeGPT Timer Momentsmall Momentbase Momentlarge Moirai-MoEsmall Moirai-MoEbase
@@
Models Moiraismall Moiraibase Moirailarge Time-MoEbase Time-MoElarge Chronossmall Chronosbase Chronoslarge TimesFM
@@
Models Moiraismall Moiraibase Moirailarge Time-MoEbase Time-MoElarge Chronossmall Chronosbase Chronoslarge
@@
Models MIRAsmall MIRAbase MIRAlarge Contiformer T-PatchGNN ODE-RNN Neural-CDE
@@
Models MIRAsmall MIRAbase MIRAlarge Moiraismall Moiraibase Moirailarge Time-MoEbase Time-MoElarge Chronossmall Chronosbase Chronoslarge

### v5

@@ 4.5 Ablation Analysis
Expert Activation Analysis.
Figure 3: Gating scores for experts across different layers in the three different frequency datasets.
We further examine the trade-off between predictive performance and inference efficiency by varying the number of activated experts ($top_{k}$), as shown in the right-hand side of Table 5. Activating two experts ($top_{2}$) achieves the best balance, reaching the lowest RMSE of 0.154 with an inference speed of 0.101 s/iter. Increasing to four experts provides no meaningful improvement in performance but increases inference time to 0.112 s/iter. Further increasing to six or eight experts leads to slower inference with diminishing returns. Conversely, using a single expert ($top_{1}$) sacrifices accuracy (0.160) for only a marginal speed gain. These findings indicate that using two experts offers an optimal trade-off between efficiency and accuracy for scalable deployment.
@@ 5 Discussion & Conclusion
Acknowledgments and Disclosure of Funding
We thank Microsoft Research for providing the computational resources that made this work possible. We are also grateful to the anonymous reviewers for their insightful comments and constructive suggestions, which significantly improved the quality of this paper. Viktor Schlegel is part of the IN-CYPHER programme and is supported by the National Research Foundation, Prime Minister’s Office, Singapore, under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. We acknowledge the support from the UMRI IDR Placement 2025 Pioneering project "Foundational AI Models for Brain Computer Interfaces and Neuro-Robotic Control" (PI: Dr. Jingyuan Sun), and The European High Performance Computing Joint Undertaking (EuroHPC JU) with project EHPC-DEV-2025D06-002 (PI: Dr. Jingyuan Sun) and project EHPC-BEN-2025B05-008 (PI: Dr. Jingyuan Sun).
@@ Appendix B Time Quantization Procedure
The procedure begins with a predefined initial resolution (e.g., 1.0) and iteratively reduces it by a shrink factor (e.g., 10) until all quantized time points become unique. In each iteration, the algorithm rounds the original timestamps to the current resolution and checks for duplicates. If uniqueness is achieved, the quantized sequence is returned. If, after a fixed number of iterations, uniqueness cannot be guaranteed, a fallback mechanism is triggered. In this case, small additive jitter is applied to break ties while preserving temporal ordering. The jitter is dynamically determined based on the magnitude of the input values, ensuring that the perturbation is both numerically safe and practically negligible. Finally, the adjusted timestamps are cast to 32-bit floating-point precision and verified again for uniqueness. If rare collisions still exist due to floating-point rounding, a second corrective jitter step is applied [90, 91].
@@ C.1 Theoretical Properties
Linear Angle Scaling.
@@
Proof. Directly from Equation (1) in the main text. The linear formulation preserves temporal causality while maintaining bounded rotation magnitudes through exponentially decaying frequencies:
$\omega_{i}=\exp\!\left(-\frac{2i}{d}\ln 10000\right),$ (14)
ensuring higher dimensions receive progressively smaller rotations.
Relative Position Encoding.
The inner product between rotated vectors depends only on their temporal difference:
$\langle R_{\Theta}(t_{i})q,\,R_{\Theta}(t_{j})k\rangle=\langle q,\,R_{\Theta}(t_{j}-t_{i})k\rangle.$ (15)
Proof. Using properties of orthogonal rotation matrices:
$R_{\Theta}(t_{i})^{\top}R_{\Theta}(t_{j})$ $=\bigoplus_{k=0}^{d/2-1}\begin{bmatrix}\cos(\omega_{k}(t_{j}-t_{i}))&\sin(\omega_{k}(t_{j}-t_{i}))\\ -\sin(\omega_{k}(t_{j}-t_{i}))&\cos(\omega_{k}(t_{j}-t_{i}))\end{bmatrix}$ (16)
$=R_{\Theta}(t_{j}-t_{i}),$ (17)
where we assume $t_{j}\geq t_{i}$; otherwise, $R_{\Theta}(t_{j}-t_{i})=R_{\Theta}(|t_{j}-t_{i}|)$ due to $\cos(-\theta)=\cos(\theta)$ and $\sin(-\theta)=-\sin(\theta)$.
C.2 Rotation Matrix Construction.
For an even-dimensional model ($d$ even), define the block-diagonal rotation operator:
$R_{\Theta}(t)=\bigoplus_{i=0}^{d/2-1}\begin{bmatrix}\cos(\omega_{i}t)&-\sin(\omega_{i}t)\\ \sin(\omega_{i}t)&\phantom{-}\cos(\omega_{i}t)\end{bmatrix}$. (18)
Key properties include:
  • Norm Preservation: $\|R_{\Theta}(t)x\|=\|x\|,\ \forall x\in\mathbb{R}^{d}$.
  • Temporal Monotonicity: $t_{1}<t_{2}\Rightarrow\theta_{i}(t_{1})<\theta_{i}(t_{2})$.
  • Differentiability: $\frac{\partial R_{\Theta}(t)}{\partial t}$ exists for all $t>0$.
@@ Appendix D Neural ODE
Gradient Computation via Adjoint Sensitivity.
For end-to-end training, gradients of a loss function $\mathcal{L}$ with respect to the ODE parameters $\theta_{\text{ODE}}$ and the initial state $h(t_{N})$ are efficiently computed using the adjoint sensitivity method [40]. This involves solving a backward-in-time ODE for the adjoint state $a(s)=\partial\mathcal{L}/\partial h(s)$:
$\frac{da(s)}{ds}=-\left(\frac{\partial f_{\text{ODE}}}{\partial h}(s,h(s);\theta_{\text{ODE}})\right)^{\top}a(s),\quad a(t_{N+1})=\frac{\partial\mathcal{L}}{\partial h^{\prime}(t_{N+1})}.$ (20)
Gradients w.r.t. parameters $\theta_{\text{ODE}}$ are then computed through integrals involving $a(s)$.
Stability Considerations.
To ensure numerical stability, the dynamics function $f_{\text{ODE}}$ is spectrally normalized, which bounds its Lipschitz constant $L$. This constrains error propagation during integration, yielding a bounded numerical error:
$\|h_{\text{ODE}}(t_{N+1})-h_{\text{true}}(t_{N+1})\|\leq\frac{M}{L}\left(e^{L\Delta t}-1\right),\quad\Delta t=t_{N+1}-t_{N},$ (21)
where $M$ represents the local truncation error of the ODE solver.
@@
0: Hidden state $h(t_{N})$, timestamps $t_{N},t_{N+1}$
0: Evolved state $\tilde{h}(t_{N+1})$
@@
2: Solve ODE: $\tilde{h}(t_{N+1})\leftarrow\texttt{odeint}(f_{\text{ODE}},h(t_{N}),[t_{N},t_{N+1}])$
3:return $\tilde{h}(t_{N+1})$
D.1 Theoretical Guarantees.
Our formulation satisfies the Picard–Lindelöf conditions for existence and uniqueness.
Theorem 1 (Existence and Boundedness).
Let $f_{\text{ODE}}$ be spectrally normalized with maximum singular value $\sigma_{\max}$, and Lipschitz continuous with constant $L=\sigma_{\max}$. Then for $\Delta t=t_{N+1}-t_{N}>0$, the state evolution admits a unique solution satisfying
$\|\tilde{h}(t_{N+1})-\tilde{h}(t_{N})\|\leq\sigma_{\max}\Delta t\,e^{\sigma_{\max}\Delta t}.$ (22)
This bounded evolution property ensures stability for long-horizon extrapolation.
@@ Appendix F Model Configurations and Training Detail
Informed by recent scaling studies [29], we design MIRA in three model scales: $MIRA_{small}$ (73M parameters), $MIRA_{base}$ (114M), and $MIRA_{large}$ (455M), enabling flexible deployment across different hardware constraints. All models are trained on up to eight NVIDIA 80GB A100 GPUs with a micro-batch size of 128 and a maximum sequence length of 512. We pre-trained one epoch with each training step processes approximately 65,000 time points. We consider forecast horizons of 24,32,48,64{24,32,48,64} for short-term and long-term evaluation. Following standard practice, we apply an auxiliary load balancing loss with weight $\alpha=0.02$ to encourage expert utilization.
@@ Appendix I Evaluation Metrics
Models
Lag-llama TimeGPT Timer Momentsmall Momentbase Momentlarge Moirai-MoEsmall Moirai-MoEbase
@@
Models
Moiraismall Moiraibase Moirailarge Time-MoEbase Time-MoElarge Chronossmall Chronosbase Chronoslarge TimesFM
@@
Models
Moiraismall Moiraibase Moirailarge Time-MoEbase Time-MoElarge Chronossmall Chronosbase Chronoslarge
@@
Models
MIRAsmall MIRAbase MIRAlarge Contiformer T-PatchGNN ODE-RNN Neural-CDE
@@
Models